# Legacy Procurement Practices Shape How U.S. Cities Govern AI: Understanding Government Employees’ Practices, Challenges, and Needs

### v2

@@ 4. Legacy Purchasing Practices Shape How Cities Govern AI
In what follows, we shine a light on the processes that municipal employees currently use to procure AI, on-the-ground. We learned that all seven interviewed cities had already started to use procured AI technologies designed for a wide set of users and goals, for example, to facilitate resident communication, increase workplace productivity, aid law enforcement, and assist bureaucratic decision-making [^3] In Appendix B, we describe each of these use cases in further detail and provide example real-world AI solutions mentioned by participants.. Notably, five out of seven cities we interviewed did not have the capacity to develop their own AI solutions internally, affirming our hypothesis that local governments rely heavily (and in many cases, entirely) on third-party AI systems.
@@ 4.1. Public procurement structures critical decisions about AI
When asked what procurement entailed, employees often described the highly structured processes their city took to award and oversee contracts with vendors. Consistent with public administration scholarship  (Curry, 2019; Halchin, 2006; for Puerto Rico, 2020), employees tended to follow a six-step process (Figure 1) for acquisitions that involved a competitive solicitation. However, not all steps occurred in every city. While some cities required employees to introduce legislation in front of City Council to approve every procurement contract, presenting an opportunity for public oversight  (Reisman et al., 2018), other cities did not need to introduce legislation for a purchase to be made. [^4] See Duchicela et al. (2023) for a more detailed review of how City Council oversight processes vary across local governments. For each step, cities applied highly standardized processes when purchasing AI. Standardization, employees explained, was important to “increase efficiency” of the procurement process and prevent government corruption (Patrucco et al., 2020; u4w, 2024). Analyzing cities’ standardized processes reveals the implicit assumptions they carry, and how existing processes both facilitate and constrain cities’ capacity to address AI harms.
Throughout interviews, we learned how cities’ solicitation processes determined if, and how, the societal impacts of AI factored into vendor selection. One common type of solicitation is a Request for Proposal (RFP), a structured process where a government outlines their needs, expectations, and desired outcomes. Interested vendors then submit detailed proposals that comprehensively address these requirements. After the solicitation has closed, cities review vendors’ proposals to select a winner. Liz, a procurement specialist, walked through her city’s RFP evaluation process by sharing her screen as she navigated her city’s procurement management software for a past IT procurement: “So all the evaluators will show up here”, Liz said, pointing to a list of ten city employees. “And when they go in, they score each vendor based on multiple criteria” such as ”Qualifications of the project team”, ”Project plan”, and others, assigning a score of 1 to 15. Once evaluators individually had assigned scores for every single proposal, they met as a committee to discuss the scores and select their “top three candidates to reach out and set up a live demonstration”.
Figure 1. Six common steps that occurred in cities’ public procurement processes. While purchases that went through a full solicitation (e.g., an RFP) proceeded linearly through these steps, as shown by the long blue arrow, AI acquired via alternative purchasing pathways such as cooperative purchasing agreements (”piggyback contracts”) or under cost thresholds (e.g., using purchasing cards) skipped past several steps, illustrated by the short blue arrows. In our paper, we discuss meaningful differences across cities’ processes. For example, not all cities had a legislation step, and some cities still required all acquisitions to go through contract negotiation. As such, this figure serves as an illustrative example of how AI procurement occurs in many (but not all) U.S. cities.
[…]
@@ 4.2. Public procurement determines who has visibility, oversight, and decision-making power over AI
Another key factor that determined the actors involved in an AI procurement was the purchasing pathway used to make the acquisition. In Figure 1, we describe the six steps that occur for procurements that go through a full solicitation. However, our interviews indicate that AI procurement often does not take this classic route, sometimes skipping competitive solicitation or contract negotiation entirely. Instead, employees leveraged a variety of alternative procurement pathways to purchase AI. In many cities, acquisitions that occurred through these alternative procurement pathways “didn’t have to go through a procurement [department]”, and thus fell outside the scope of existing accountability measures for government purchasing. Due to broader shifts in the AI landscape, namely the availability of low- and no-cost AI tools, many AI acquisitions did not involve a competitive solicitation because they were under specified cost thresholds (Duchicela et al., 2023) that would require going to solicitation. Local policies specified that municipal employees could make purchases under a certain dollar amount (which varied city-to-city) at their own discretion, using a government-issued purchasing card. Cost threshold policies, procurement director Olivia explained, were established to “triage risk, by defining the [riskiest projects] based on dollar value”. Types of AI tools that fell under cost thresholds included free online services like ChatGPT, services with paid subscription models, or AI donated through academic collaborations, foundations, or from for-profit companies [^5] For instance, Palantir, a for-profit company, donated its predictive policing technology as a “philanthropic gift” to be used by the City of New Orleans, free of charge  (Stanley, 2018)..
The governance processes applicable to procurements under cost thresholds varied significantly from city to city. Adrian, an IT leader, reflected on how technology acquisitions under cost thresholds were particularly difficult for their department to even be aware of until after they were already purchased: “[For purchases] below \$x0,000, there’s few oversight of regulatory mechanisms to control, or even have visibility of what departments do,” they explained. “It’s not routed through a centralized control mechanism.” As a result, individual employees could purchase and deploy free and low-cost AI tools (e.g., with prices under \$x00 per month) at their own discretion, without notifying others in the city. In contrast, Mari’s city still routed ”anything involving a company or vendor” through a centralized procurement review: “Even if it’s an in kind donation, it is still tied to a scope-of-work and a contract.”
@@ 4.3. Public procurement shapes the socio-technical values of AI assessment
Public procurement norms, particularly cost-effectiveness, played an important role in how employees assessed AI proposals. “When we make decisions, we think about how we’re stewards of residents’ tax money,” explained Emma, an IT procurement specialist. Indeed, cost was central factor that influenced employees’ AI purchasing decisions. Employees shared many anecdotes where they ultimately decide not to procure an AI system because it was too expensive [P4, P6, P8, P14]. While some employees like Hugo expressed enthusiasm about the increased availability of low-cost consumer AI tools (such as a generative AI pilot that only cost Hugo’s city \$5), others expressed concern over the “extractive capabilities” and hidden costs of free AI tools. Adrian recalled an incident where an employee used a transcription AI tool that did not have “a consensual model for data collection”. Although the transcription tool was free, its use could result in “divulging resident information” or “more secure” government business, data that the vendor could use to improve its models. Adrian’s story highlights a tension between the long-standing norm of cost-effectiveness and the new reality of unique risks posed by low- and no-cost AI.
@@
One significant difference between cities’ IT review processes was the degree to which they prioritized values such as data privacy and protection from government surveillance. Although these considerations were never mentioned in most employees’ descriptions of their review criteria, participants in a small number of cities called them out as organizational priorities. Understanding these priorities, we learned, was critical to understanding how cities procured AI. For example, cities with established privacy programs often tasked privacy personnel to work on AI-focused initiatives, and adapted existing privacy governance infrastructure to include AI-specific considerations. Similarly, cities with surveillance ordinances had already established governance processes such as community oversight mechanisms for qualifying AI technologies  (Young et al., 2019). Organizational values also shaped the way employees conceptualized AI’s benefits and risks: Kai shared that it was the ”privacy lens” adopted by her organization that led her to ask vendors questions about training dataset curation and consent. In contrast, employees in cities without surveillance or privacy policies spoke considerably more positively about AI-powered “smart city”  (Whitney et al., 2021) or policing technologies. Thus, cities’ (lack of) existing privacy or surveillance policies played a critical role in determining what risks were considered, what mitigations were requested, and what technologies were ultimately purchased and deployed.
@@ 4.4. Emerging practices: New actions cities have taken to govern AI
Intervening within existing procurement processes. The most common action considered by cities was to modify their existing contract and solicitation templates to include additional items for AI-related procurements. Cities instituted additional reporting requirements for AI systems in their technology RFPs, adopting language from the GovAI Coalition’s AI FactSheet  (gov, 2024a) of questions for vendors. The factsheet asks vendors to report “essential technical details” such as on what data the AI was trained, under what conditions the system was tested, the values of relevant performance metrics, and steps that vendors have taken to promote values of fairness, robustness, and explainability. Cities also adapted their contracting practices by developing additional contract language specific to AI procurements. Many employees adopted language from GovAI’s Vendor Agreement  (gov, 2024b), a “plug-and-play legal addendum” of requirements for vendors, e.g., to develop an AI incident response plan, remediate AI incidents, and provide governments with a means to monitor and audit AI performance. Most interviewed cities did not plan to adopt GovAI’s resource templates as-is, but instead chose to “selectively adopt” the items that felt most important to their city.
@@
Establishing an AI review process. Three interviewed cities established an internal AI review process overseen by IT employees, that occurred outside of the formal review processes of procurement (e.g., scorecards). When a city employee wanted to acquire an AI solution, they initiated an AI review by contacting these IT specialists directly, e.g., by filling out a form or “opening a ticket”. IT employees or committees trained to assess AI systems then reviewed the employees’ request. AI reviews often involved conducting risk assessments to help city employees understand potential positive and negative impacts of the system. Employees used pre-acquisition risk assessments to shape subsequent conversations with AI vendors, for instance, by requesting specific risk mitigation steps to be implemented in the AI system’s design. Some cities used risk assessments to institute additional governance steps for “high-risk” systems, such as instituting internal usage protocols, public reporting requirements, and regular post-deployment monitoring. We discuss governments’ emerging AI risk assessment practices in detail in Appendix C.
@@ 5. Three Key Challenges at the Heart of AI Procurement Reform
Although some employees pointed us to success stories where their purchasing processes enabled them to walk away from red flags or negotiate for mitigations, we found that cities’ present AI procurement practices often fell short of their ambition to protect the public from risks and harms of AI. In what follows, we reflect on three challenges that often motivate, but are not adequately addressed by existing AI procurement reform efforts. While considerable progress has been made, we believe that more work is urgently needed to (1) address information asymmetries between governments and vendors, (2) support cities in asking more of AI vendors, and (3) support cities in sharing and assuming ongoing responsibilities of AI governance. Under each challenge, we identify opportunities for how the FAccT community might contribute towards addressing local governments’ needs for support.
@@ 5.1. Information asymmetry and (un)preparedness for AI
Employees shared a variety of methods they used to increase their organization’s AI literacy and preparedness, such as recruiting employees with AI expertise, resource sharing efforts (e.g., through GovAI), and participating in intergovernmental AI task forces  (Dwyer, 2025) or grassroots peer networks. Some employees began collaborations with academic researchers, for example, to evaluate a pilot of a generative AI productivity tool. Finally, some cities hired paid consultants (e.g., from Gartner) to provide strategic advising (e.g., conduct market research) on AI. Employees in smaller cities were less likely to be aware of existing resources or peer networks to learn about AI, and felt overwhelmed by the broader information ecosystem of “hype cycles” surrounding emerging generative AI technologies.
@@
Recommendations for future work. Although much progress has been made in developing resources to support governments’ AI procurement practices, our findings reveal that there is still an urgent need for government education and support. Misconceptions surrounding AI capabilities, often perpetuated by AI vendors, resulted in power imbalances where cities’ existing review criteria for software proved insufficient to anticipate AI harms. We believe that the FAccT community has a critical role to play in doing this translation work. Our conversations revealed a number of avenues through which members of the public can influence AI policymaking, such as by collaborating on research  (Stapleton et al., 2022b), joining peer networks such as the GovAI Coalition  (gov, 2024c), or even simply reaching out to one’s own government to understand their AI governance practices and needs for support. Following interviews, our research team curated and shared educational resources that we hoped would address employees’ open questions about AI. These conversations, we learned, were often a precipitating event that led interviewed cities to join AI-focused peer networks, revise their procurement practices, and even establish new ethical AI governance policies. The FAccT conference can explore models to build community and develop trust-based relationships across scholars, community advocates, and government employees, e.g., following models such as the Public Tech Leadership Collaborative  (ds_, 2025).
@@
5.2. Cities lack leverage in relationships with AI vendors
Several employees from cities with new AI governance measures described positive experiences with ”good vendors” who were transparent and responsive. For instance, an AI vendor’s openness to disclose system limitations in their autonomous drone technology helped a city refine its usage policies to avoid its deployment in low-accuracy environments. Despite these successes, we heard many more stories of uncooperative AI vendors that failed to provide basic information about their AI system, amend their contracting terms, or implement mitigations requested by the city. Interviewed employees repeatedly felt that they lacked leverage in advocating on behalf of their city.
AI vendors frequently withheld critical information about their AI systems from governments, hindering employees’ ability to assess risks. Vendors often invoked intellectual property claims to deny basic requests, such as responding to AI RFP items or completing AI FactSheets. Consistent with past research  (Brauneis and Goodman, 2018; Sloane et al., 2021; Mulligan and Bamberger, 2019; Kawakami et al., 2024a), participants were often denied access to vendors’ AI models (e.g., model weights) and training datasets. Additional denied requests included information about the presence of copyrighted content in training data [P3, P5], whether city data would be used to train vendor’s models [P1, P17], and disaggregated accuracy measures, e.g., across demographic groups [P2, P7, P10].
While bad vendor behavior could serve as an informative signal when making purchasing decisions, relational challenges still arose for procurements that were already under-contract. Past procurements were important to consider given the prevalence of vendors’ “scope creep”  (Stevens, 2022) that introduced AI under existing contracts. Also, many cities’ AI review processes were so new that they did not apply for most AI they already owned. Employees shared anecdotes where AI vendors did not provide adequate personnel training [P4, P17], refused to give cities access to data collected on city employees or residents [P2, P14, P15], and refused to let cities opt out of new AI features [P1, P17] or their data being used for AI training [P5, P13, P16]. In some cases, cities responded by “shutting off” purchased services, such as a city that stopped paying for Zoom because they were “uncomfortable with their AI notetaker”. Yet, several employees discussed past procurements where they struggled to break ties with vendors. For instance, employees shared stories where they struggled to disable AI features that were deeply embedded in necessary technical infrastructure, such as new AI productivity features introduced into enterprise software [P1, P3, P5, P6, P11, P15].
@@ 5.2. Cities lack leverage in relationships with AI vendors
Recommendations for future work. We found that many AI vendors refused to voluntarily comply with cities’ requests to provide basic information or implement simple harm mitigation steps. This context of low compliance is critical to consider in this moment of unprecedented attention targeting governments’ AI procurement practices. Several interviewed employees expressed their frustration in having to advocate for basic ethical behaviors (e.g., transparency about how the vendors would use resident data) that they felt should be minimum requirements to even be able to participate in the market (e.g., enforced as consumer protection rights or federal data privacy legislation). These stories call into question the hope implicit to initiatives that focus primarily on governments’ power as purchasers: that when it comes to protecting the public interest, the market of public sector AI vendors will simply regulate itself.
When vendors failed to meet expectations, a variety of factors such as contract agreements or conflicting incentives across government employees meant that procurements still moved forward. Future research efforts can more deeply interrogate the role that legacy purchasing practices like contracts play in constraining cities’ negotiation power, and ways that cities might adopt their processes to address these issues head-on (e.g., by explicitly giving AI reviewers the power to reject a proposed AI system due to its potential to cause harm).

### v3

@@ 4. Legacy Purchasing Practices Shape How Cities Govern AI
Figure 1. Six common steps that occurred in cities’ public procurement processes. While purchases that went through a full solicitation (e.g., an RFP) proceeded linearly through these steps, as shown by the long blue arrow, AI acquired via alternative purchasing pathways such as cooperative purchasing agreements (”piggyback contracts”) or under cost thresholds (e.g., using purchasing cards) skipped past several steps, illustrated by the short blue arrows. In our paper, we discuss meaningful differences across cities’ processes.
[…]
In what follows, we shine a light on the processes that municipal employees currently use to procure AI, on-the-ground. We learned that all seven interviewed cities had already started to use procured AI technologies designed for a wide set of users and goals, for example, to facilitate resident communication, increase workplace productivity, aid law enforcement, and assist bureaucratic decision-making. [^4] In Appendix B, we describe each of these use cases in further detail and provide example real-world AI solutions mentioned by participants. Notably, five out of seven cities we interviewed did not have the capacity to develop their own AI solutions internally, affirming our hypothesis that local governments rely heavily (and in many cases, entirely) on third-party AI systems.
@@ 4.1. Public procurement structures critical decisions about AI
When asked what procurement entailed, employees often described the highly structured processes their city took to award and oversee contracts with vendors. Consistent with public administration scholarship  (Curry, 2019; Halchin, 2006; Financial Oversight & Management Board for Puerto Rico, 2020), employees tended to follow a six-step process (Figure 1) for acquisitions that involved a competitive solicitation. However, not all steps occurred in every city. While some cities required employees to introduce legislation in front of City Council to approve every procurement contract, presenting an opportunity for public oversight  (Reisman et al., 2018), other cities did not need to introduce legislation for a purchase to be made. [^5] See Duchicela et al. (2023) for a more detailed review of how City Council oversight processes vary across local governments. For each step, cities applied highly standardized processes when purchasing AI. Standardization, employees explained, was important to “increase efficiency” of the procurement process and prevent government corruption (Patrucco et al., 2020; U4 Anti-Corruption Resource Centre, 2024). Analyzing cities’ standardized processes reveals the implicit assumptions they carry, and how existing processes both facilitate and constrain cities’ capacity to address AI harms.
Throughout interviews, we learned how cities’ solicitation processes determined if, and how, the societal impacts of AI factored into vendor selection. One common type of solicitation is a Request for Proposal (RFP), a structured process where a government outlines their needs, expectations, and desired outcomes. Interested vendors then submit detailed proposals that comprehensively address these requirements. After the solicitation has closed, cities review vendors’ proposals to select a winner. Liz, a procurement specialist, walked through her city’s RFP evaluation process by sharing her screen as she navigated her city’s procurement management software for a past IT procurement: “So all the evaluators will show up here”, Liz said, pointing to a list of ten city employees. “And when they go in, they score each vendor based on multiple criteria” such as ”Qualifications of the project team”, ”Project plan”, and others, assigning a score of 1 to 15. Once evaluators individually had assigned scores for every proposal, they met as a committee to discuss the scores and select their “top three candidates to reach out and set up a live demonstration”.
@@ 4.2. Public procurement determines who has visibility, oversight, and decision-making power over AI
Another key factor that determined the actors involved in an AI procurement was the purchasing pathway used to make the acquisition. In Figure 1, we describe the six steps that occur for procurements that go through a full solicitation. However, our interviews indicate that AI procurement often does not take this classic route, sometimes skipping competitive solicitation or contract negotiation entirely. Instead, employees leveraged a variety of alternative procurement pathways to purchase AI. In many cities, acquisitions that occurred through these alternative procurement pathways “didn’t have to go through a procurement [department]”, and thus fell outside the scope of existing accountability measures for government purchasing. Due to broader shifts in the AI landscape, namely the availability of low- and no-cost AI tools, many AI acquisitions did not involve a competitive solicitation because they were under specified cost thresholds (Duchicela et al., 2023) that would require going to solicitation. Local policies specified that municipal employees could make purchases under a certain dollar amount (which varied city-to-city) at their own discretion, using a government-issued purchasing card. Cost threshold policies, procurement director Olivia explained, were established to “triage risk, by defining the [riskiest projects] based on dollar value”. Types of AI tools that fell under cost thresholds included free online services like ChatGPT, services with paid subscription models, or AI donated through academic collaborations, foundations, or from for-profit companies. [^6] For example, Palantir, a for-profit company, donated its predictive policing technology as a “philanthropic gift” to be used by the City of New Orleans, free of charge  (Stanley, 2018).
The governance processes applicable to procurements under cost thresholds varied significantly from city to city. Adrian, an IT leader, reflected on how technology acquisitions under cost thresholds were particularly difficult for their department to even be aware of until after they were already purchased: “[For purchases] below \$x0,000, there’s few oversight of regulatory mechanisms to control, or even have visibility of what departments do,” Adrian explained. “It’s not routed through a centralized control mechanism.” As a result, individual employees could purchase and deploy free and low-cost AI tools (e.g., with prices under \$x00 per month) at their own discretion, without notifying others in the city. In contrast, Mari’s city still routed ”anything involving a company or vendor” through a centralized procurement review: “Even if it’s an in kind donation, it is still tied to a scope-of-work and a contract.”
@@ 4.3. Public procurement shapes the socio-technical values of AI assessment
Public procurement norms, particularly cost-effectiveness, played an important role in how employees assessed AI proposals. “When we make decisions, we think about how we’re stewards of residents’ tax money,” explained Emma, an IT procurement specialist. Indeed, cost was a central factor that influenced employees’ AI purchasing decisions. Employees shared many anecdotes where they ultimately decide not to procure an AI system because it was too expensive [P4, P6, P8, P14].
On the other hand, some AI technologies’ low price tag, or promised cost-saving capabilities, served as an incentive to some. While employees like Hugo expressed enthusiasm about the increased availability of low-cost consumer AI tools (such as a generative AI pilot that only cost Hugo’s city \$5), others expressed concern over the “extractive capabilities” and hidden costs of free AI tools. Adrian recalled an incident where an employee used a transcription AI tool that did not have “a consensual model for data collection”. Although the transcription tool was free, its use could result in “divulging resident information” or “more secure” government business, data that the vendor could use to improve its models. Adrian’s story highlights a tension between the long-standing norm of cost-effectiveness and the new reality of unique risks posed by low- and no-cost AI.
@@
One significant difference between cities’ IT review processes was the degree to which they prioritized values such as data privacy and protection from government surveillance. Although these considerations were never mentioned in most employees’ descriptions of their review criteria, participants in a small number of cities called them out as organizational priorities. Understanding these priorities, we learned, was critical to understanding how cities procured AI. For example, cities with established privacy programs often tasked privacy personnel to work on AI-focused initiatives, and adapted existing privacy governance infrastructure to include AI-specific considerations. Similarly, cities with surveillance ordinances had already established governance processes such as community oversight mechanisms for qualifying AI technologies  (Young et al., 2019).
Organizational values also shaped the way employees conceptualized AI’s benefits and risks: Kai shared that it was the ”privacy lens” adopted by her organization that led her to ask vendors questions about training dataset curation and consent. In contrast, employees in cities without surveillance or privacy policies spoke considerably more positively about AI-powered “smart city”  (Whitney et al., 2021) or policing technologies. Thus, cities’ (lack of) existing privacy or surveillance policies played a critical role in determining what risks were considered, what mitigations were requested, and what technologies were ultimately purchased and deployed.
@@ 4.4. Emerging practices: New actions cities have taken to govern AI
Intervening within existing procurement processes. The most common action considered by cities was to modify their existing contract and solicitation templates to include additional items for AI-related procurements. Cities instituted additional reporting requirements for AI systems in their technology RFPs, adopting language from the GovAI Coalition’s AI FactSheet  (The Government AI Coalition, 2024a) of questions for vendors. The factsheet asks vendors to report “essential technical details” such as on what data the AI was trained, under what conditions the system was tested, the values of relevant performance metrics, and steps that vendors have taken to promote values of fairness, robustness, and explainability. Cities also adapted their contracting practices by developing additional contract language specific to AI procurements. Many employees adopted language from GovAI’s Vendor Agreement  (The Government AI Coalition, 2024b), a “plug-and-play legal addendum” of requirements for vendors, e.g., to develop an AI incident response plan, remediate AI incidents, and provide governments with a means to monitor and audit AI performance. Most interviewed cities did not plan to adopt GovAI’s resource templates as-is, but instead chose to “selectively adopt” the items that felt most important to their city.
@@
Establishing an AI review process. Three interviewed cities established an internal AI review process overseen by IT employees, that occurred outside of the formal review processes of procurement (e.g., scorecards). When a city employee wanted to acquire an AI solution, they initiated an AI review by contacting these IT specialists directly, e.g., by filling out a form or “opening a ticket”. IT employees or committees trained to assess AI systems then reviewed the employees’ request. AI reviews often involved conducting risk assessments to help city employees understand potential impacts of the system. Employees used pre-acquisition risk assessments to shape subsequent conversations with AI vendors, for instance, by requesting specific risk mitigation steps to be implemented in the AI system’s design. Some cities used risk assessments to institute additional governance steps for “high-risk” systems, such as instituting internal usage protocols, public reporting requirements, and regular post-deployment monitoring. We discuss governments’ emerging AI risk assessment practices in detail in Appendix D.
@@ 5. Three Key Challenges at the Heart of AI Procurement Reform
Although some employees pointed us to success stories where their purchasing processes enabled them to walk away from red flags or negotiate for mitigations, we found that cities’ present AI procurement practices often fell short of their ambition to protect the public from risks and harms of AI. In what follows, we reflect on three challenges that often motivate, but are not adequately addressed by existing AI procurement reform efforts. We revisit challenges previously surfaced by past scholarship  (Veale et al., 2018; Engstrom and Ho, 2021; Autio et al., 2023; Kawakami et al., 2024b; Hickok, 2024) to reflect on the considerable progress that has been made, and the work that is still urgently needed to (1) address information asymmetries between governments and vendors, (2) support cities in asking more of AI vendors, and (3) support cities in sharing ongoing responsibilities of AI governance. Under each challenge, we identify opportunities for how the FAccT community might contribute towards addressing local governments’ needs for support.
@@ 5.1. Information asymmetry and (un)preparedness for AI
Employees shared a variety of methods they used to increase their organization’s AI literacy and preparedness, such as recruiting employees with AI expertise, resource sharing efforts (e.g., through GovAI), and participating in intergovernmental AI task forces  (Dwyer, 2025) or grassroots peer networks. Some employees began collaborations with academic researchers, for example, to evaluate a pilot of a generative AI productivity tool. Finally, some cities hired paid consultants (e.g., from Gartner) to provide strategic advising (e.g., conduct market research) on AI. Employees in smaller cities were less likely to be aware of existing resources or peer networks to learn about AI, and felt overwhelmed by the ecosystem of “hype cycles” surrounding emerging generative AI technologies.
@@
Recommendations for future work
Although much progress has been made in developing resources to support governments’ AI procurement practices, our findings reveal that there is still an urgent need for government education and support. Misconceptions surrounding AI capabilities, often perpetuated by AI vendors, resulted in power imbalances where cities’ existing review criteria for software proved insufficient to anticipate AI harms. We believe that the FAccT community has a critical role to play in doing this translation work. Our conversations revealed a number of avenues through which members of the public can influence AI policymaking, such as by collaborating on research  (Stapleton et al., 2022b), joining peer networks such as the GovAI Coalition  (The Government AI Coalition, 2024c), or even simply reaching out to one’s own government to understand their AI governance practices and needs for support. Following interviews, our research team curated and shared educational resources that we hoped would address employees’ open questions about AI. These conversations, we learned, were often a precipitating event that led interviewed cities to join AI-focused peer networks, revise their procurement practices, and even establish new ethical AI governance policies. The FAccT conference can explore models to build community and develop trust-based relationships across scholars, community advocates, and government employees, e.g., following models such as the Public Tech Leadership Collaborative  (Data & Society, 2025).
@@
5.2. Challenge: Cities lack leverage in relationships with AI vendors.
Several employees from cities with new AI governance measures described positive experiences with ”good vendors” who were transparent and responsive. For instance, an AI vendor’s openness to disclose system limitations in their autonomous drone technology helped a city refine its usage policies to avoid its deployment in low-accuracy environments. Despite these successes, we heard many more stories of uncooperative AI vendors that failed to provide basic information about their AI system, amend their contracting terms, or implement mitigations requested by the city. We found that employees repeatedly felt that they lacked leverage in advocating on behalf of their city throughout the procurement process.
AI vendors frequently withheld critical information about their AI systems from governments, hindering employees’ ability to assess risks. Vendors often invoked intellectual property claims to deny basic requests, such as responding to AI RFP items or completing AI FactSheets. Consistent with past research  (Brauneis and Goodman, 2018; Sloane et al., 2021; Mulligan and Bamberger, 2019; Kawakami et al., 2024a), participants were often denied access to vendors’ AI models (e.g., model weights) and training datasets. Additional denied requests included information about the presence of copyrighted content in training data [P3, P5], whether city data would be used to train vendor’s models [P1, P17], and asks for disaggregated accuracy measures, e.g., across demographic groups [P2, P7, P10].
While bad vendor behavior could serve as an informative signal when making purchasing decisions, relational challenges still arose for procurements that were already under-contract. Past procurements were important to consider given the prevalence of vendors’ “scope creep”  (Stevens, 2022) that introduced AI under existing contracts. Employees shared anecdotes where AI vendors did not provide adequate personnel training [P4, P17], refused to give cities access to data collected on city employees or residents [P2, P14, P15], and refused to let cities opt out of new AI features [P1, P17] or their data being used for AI training [P5, P13, P16]. In some cases, cities responded by “shutting off” purchased services, such as a city that stopped paying for Zoom because they were “uncomfortable with their AI notetaker.” Yet, several employees discussed past procurements where they struggled to break ties with vendors. For instance, employees shared stories where they struggled to disable AI features that were deeply embedded in necessary technical infrastructure, such as new AI productivity features introduced into enterprise software [P1, P3, P5, P6, P11, P15].
@@
Recommendations for future work
We found that many AI vendors refused to voluntarily comply with cities’ requests to provide basic information or implement simple harm mitigation steps. This context of low compliance is critical to consider in this moment of unprecedented attention targeting governments’ AI procurement practices. Several interviewed employees expressed their frustration in having to advocate for basic ethical behaviors (e.g., transparency about how the vendors would use resident data) that they felt should be minimum requirements to even be able to participate in the market (e.g., enforced as consumer protection rights or through federal data privacy legislation). These stories call into question the hope implicit to initiatives that focus primarily on governments’ power as purchasers: that when it comes to protecting the public interest, the market of public sector AI vendors will simply regulate itself.
When vendors failed to meet expectations, a variety of factors such as contract agreements or conflicting incentives across employees meant that procurements still moved forward. Future research efforts can more deeply interrogate the role that legacy purchasing practices play in constraining cities’ negotiation power, and ways that cities might adapt their processes to address these issues head-on, e.g., by explicitly giving AI reviewers the power to reject a proposed AI system due to its potential to cause harm.